OpenStudy (anonymous):

find the slope-intercept form for the line satisfying the following conditions. x-intercept 2,y-intercept 1/3 the slope-intercept form for the line is_____

OpenStudy (anonymous):

@Atkinsoha

OpenStudy (anonymous):

Okay, so what is the slope intercept form?

OpenStudy (anonymous):

c=1/3 point( 2;0) y=mx+c 0=m(2)+1/3 m=-1/6 y=-1/6X +1/3

OpenStudy (anonymous):

@physicsboffin how you do tht

OpenStudy (anonymous):

@Atkinsoha x-intercept 2, y-intercept 1/3

OpenStudy (anonymous):

okay, so slope intercept form is y=mx+b

OpenStudy (anonymous):

so b is the y intercept. so we know that y=mx+ 1/3

OpenStudy (anonymous):

they say that x-inertcept is 2...this implies that (2;0)...when x=2 y=o this gives you 1 point...then u have your y-intercept which is c in the equation Y=mX+c

OpenStudy (anonymous):

now we find the slope. I usually make a quick sketch to get an approximation of the x and y intercepts. We know they are both on the line, so we do slope=rise/run with is (1/3) / 2 Which is the same as (1/3)(2/1) = 1/6

OpenStudy (anonymous):

so now we have y=(1/6)x + (1/3)

OpenStudy (anonymous):

Now graph it, and we notice that the graph I've given isn't correct. Sometimes this happens. Let's try y=-(1/6)x + (1/3)

OpenStudy (anonymous):

When dealing with the slope, it can go positive or negative, and sometimes I confuse myself with which one it is, so be sure to graph, and try both ways to ensure the correct answer. So your slope intercept form is: y= (-1/6)x + (1/3)

OpenStudy (anonymous):

thanks again